OpenStudy (anonymous):

find derivative of a function f(x)=11-4x^3

zepdrix (zepdrix):

Have you learned the `Power Rule for Derivatives` yet? Or do you need to use the `Limit Definition of the Derivative` in order to solve this? :)

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

yes to which? :) lol

OpenStudy (anonymous):

power rule 4 derivatives

zepdrix (zepdrix):

Ok cool c:

OpenStudy (anonymous):

like x^3= 3x^2? is that d power rule?

zepdrix (zepdrix):

yes good c:

OpenStudy (anonymous):

hahha. im not sure.

zepdrix (zepdrix):

\[\large f(x)=11-4x^3\] See how the first term contains no variable? There is no variation, no change in 11. It is always constant. The derivative measures the rate of change. So how much does a constant change? Zero amount! It doesn't change, it's always 11 in this case. That's a long-winded explanation of why the derivative of a constant is 0. On the second term, we'll apply the power rule for derivatives (in this order): ~Bring the power down as a coefficient on the x. ~Decrease the power by 1.

zepdrix (zepdrix):

\[\large -4x^3 \qquad \rightarrow \qquad -4\cdot3x^{3-2}\]Understand what we did there? We brought the 3 down as multiplication, and then we decrease the power by 1.

OpenStudy (anonymous):

okay.. i got it. and then

OpenStudy (anonymous):

-12x^2

zepdrix (zepdrix):

\[\large f(x)=11-4x^3\]\[\large f'(x)=0-4\cdot3x^2\]We place a prime on the f, to show that we've taken it's derivative. I see you have simplified it, good job! :)\[\large f'(x)=-12x^2\]

OpenStudy (anonymous):

hold on. i have another one. it has square root

OpenStudy (anonymous):



zepdrix (zepdrix):

So this is what the problem looks like? \[\large 2\sqrt x-\frac{3}{x^2}\]

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

We can rewrite the square root term as a `Rational Expression`. A Rational Expression is one which allows us to write the powers AND roots as exponents using fraction notation.\[\large \sqrt x=x^{1/2}\]The 2 in the denominator is to let us know that it's a 2nd root, or "square" root. So that allows us to write our problem as,\[\large 2x^{1/2}-\frac{3}{x^2}\]

zepdrix (zepdrix):

To deal with the second term, we'll have to remember some rules of exponents. We can bring the x up to the numerator, by applying a negative to the power.\[\large 2x^{1/2}-3x^{-2}\]See how the 2 changed to a -2?

zepdrix (zepdrix):

From here, everything is in a really nice format, so we can apply the `Power Rule` once again.

OpenStudy (anonymous):

yes ;)

OpenStudy (anonymous):